Project

General

Profile

Actions

Physics shift 19 07 2015

Physics Shift Report

Start Date: 19/07/2015
End Date: 23/07/2015
Start Run: 7256
End Run: 7290
Physics shifter: John Nugent

Summary

Continue cool down of magnets and take tracker data at half fields for tracker debug and alignment. TOF calibrations were also performed throughout.

Run documentation

Problems with the MICE DAQ were reported for both the EMR and the tracker. Despite data being take on the night of the 21/07/2015 the data was not analysable.

Elog

Log entries for all shifts complete with descriptions of run purpose included.

Run Condition Summary Spreadsheet/CDB

Target depth in CDB and spreadsheet do not agree for runs taken in this period. Otherwise spreadsheet and CDB completed and in agreement.

Analysis

https://micewww.pp.rl.ac.uk/projects/analysis/wiki/Rogers_analysis_2015-07-24

A summary of the first pass at the data analysis can be found at the above link. The link shows the residuals for the reconstructed kappa & theta which is the angle between the solenoid and the tracker. And the first helical trackers in field from the tracker.

Updated by Nugent, John over 8 years ago ยท 1 revisions